### v1

# Benchmarking Large Language Models for Cyberbullying Detection in Real-World YouTube Comments

Abstract
As online platforms grow, comment sections increasingly host harassment that undermines user experience and well-being. This study benchmarks three leading large language models, OpenAI GPT-4.1, Google Gemini 1.5 Pro, and Anthropic Claude 3 Opus, on a corpus of 5 080 YouTube comments sampled from high-abuse threads in gaming, lifestyle, food vlog, and music channels. The dataset comprises 1 334 harmful and 3 746 non-harmful messages in English, Arabic, and Indonesian, annotated independently by two reviewers with substantial agreement (Cohen’s $\kappa$ = 0.83). Using a unified prompt and deterministic settings, GPT-4.1 achieved the best overall balance with an F1 score of 0.863, precision of 0.887, and recall of 0.841. Gemini flagged the highest share of harmful posts (recall = 0.875) but its precision fell to 0.767 due to frequent false positives. Claude delivered the highest precision at 0.920 and the lowest false-positive rate of 0.022, yet its recall dropped to 0.720. Qualitative analysis showed that all three models struggle with sarcasm, coded insults, and mixed-language slang. These results underscore the need for moderation pipelines that combine complementary models, incorporate conversational context, and fine-tune for under-represented languages and implicit abuse. A de-identified version of the dataset and full prompts is publicly released to promote reproducibility and further progress in automated content moderation.
Index Terms:
Artificial intelligence, content moderation, cyberbullying, hate speech detection, large language models, natural language processing
I Introduction
Online platforms now mediate much of everyday communication, with more than 4.95 billion active social-media accounts worldwide in 2023 [1]. That ubiquity carries risk: recent studies report that between 6.5% and 35.4% of adolescents in the U.S. and Europe have experienced cyberbullying [2], and longitudinal studies link such exposure to elevated anxiety, depression, and suicidal ideation [3]. As the scale and psychological impact of online abuse grow, so does the urgency of developing reliable, context-aware moderation methods, especially for platforms popular with adolescents such as YouTube and TikTok.
I-A Background
Cyberbullying has become a major digital-safety concern because abusive material is persistent, anonymous, and can spread instantaneously across networks [4]. As global connectivity and social-media adoption have expanded, online harassment has become more visible, with increasing concern around its scale and severity [5].
Comment sections on social networks, video-sharing sites, and forums are particular hotspots. In a 2024 U.S. survey of parents, 79 % said their children had encountered cyberbullying on YouTube [6]. A cross-national analysis of 180 000 adolescents in 42 countries also linked problematic social-media use to both cyberbullying victimisation and perpetration, with stronger effects among girls [7].
In response, platforms increasingly deploy artificial-intelligence (AI) systems to moderate user-generated content at scale [8, 9]. Understanding where current large language models (LLMs) succeed and fail is therefore critical for building safer online spaces.
I-B Motivation and Problem Statement
AI systems are now central to content moderation, especially on large platforms where the volume of user-generated posts exceeds human capacity [10]. Yet despite advances in natural language processing, enforcement remains inconsistent. Abusive comments that rely on sarcasm, coded language, or emotional manipulation often slip through AI moderation, even when they clearly violate platform policies. These subtleties continue to pose challenges for automated systems, which often misread tone or miss indirect expressions of harm [11].
This is especially dangerous in cyberbullying, where the impact of a single overlooked comment can be deeply personal [12]. Unlike hate speech or spam, bullying is often subtle: expressed through mockery, group pressure, or repeated jabs that may seem harmless in isolation but accumulate over time. Many academic benchmarks fail to capture this complexity, relying on synthetic or crowd-sourced datasets with clear-cut abuse [13].
This study addresses that gap by evaluating three state-of-the-art LLMs, GPT-4, Gemini, and Claude, on authentic YouTube comments from videos with documented cyberbullying. By focusing on nuanced, emotionally loaded content, we examine not just whether models detect harmful language, but how well they handle ambiguity, cultural context, and tone.
I-C Objectives of the Study
The work pursues four specific objectives:
  1. 1.
    Benchmark performance. Compare OpenAI GPT, Google Gemini, and Anthropic Claude in detecting cyberbullying within real YouTube data.
  2. 2.
    Quantify error patterns. Measure false-positive and false-negative rates and identify recurring misclassification themes.
  3. 3.
    Conduct qualitative error analysis. Manually inspect model outputs to assess contextual understanding and pinpoint linguistic edge cases.
  4. 4.
    Assess robustness across nuance. Evaluate how well each model handles abusive content that is linguistically ambiguous, culturally specific, or indirectly expressed.
II Related Work
II-A AI Content Moderation and Toxic Language Detection
Automated content moderation has evolved from simple keyword-based filters to machine learning classifiers and, more recently, transformer-based models. Early systems relied on blacklists or logistic regression trained on hand-labeled corpora, but struggled with negation, sarcasm, and informal language. The advent of BERT [14], RoBERTa [15], and their derivatives enabled context-aware classification by capturing bidirectional dependencies and subtle phrasing nuances. These models improved performance on public benchmarks such as the Jigsaw Toxic Comment dataset and HateXplain [16], driving adoption in industry moderation pipelines.
However, most benchmarks are composed of isolated sentences and short comments labeled through crowdwork. They rarely reflect the conversational, multi-turn, or emotionally charged dynamics seen in real-world platforms. Furthermore, these datasets often overrepresent overt toxicity while underrepresenting gray-area content such as mockery, dismissive sarcasm, or personal digs. As a result, models trained on them perform well in benchmark settings but struggle in live deployments where language is more ambiguous and emotionally expressive [17, 18].
II-B Challenges in Detecting Implicit and Nuanced Abuse
Detecting implicit harm remains one of the most difficult challenges in AI-based moderation. Sarcasm, coded language, cultural slang, and subtextual insults frequently evade model detection, even in large LLMs. For example, a phrase like “She’s always such a queen, right?” may be an insult depending on tone and context, but appears harmless in isolation. Work on sarcasm detection [19], adversarial misspellings, irony detection, and tone modeling has attempted to address this gap, but progress remains limited [20].
Recent research also suggests that models often overflag emotionally intense or sensitive language even when used in supportive contexts, leading to user frustration and moderation fatigue. This tension between underflagging subtle harm and overflagging emotional expression makes fine-tuned moderation especially difficult. Studies have proposed integrating tone analysis, conversational history [22], and speaker intent as ways to improve subtle abuse detection, but these are rarely deployed at scale.
II-C Moderating Multilingual and Underresourced Content
While English dominates most training corpora, abuse occurs in every language, and much of it is multilingual, transliterated, or mixed with emojis and slang. Studies show a consistent drop in model accuracy when applied to languages such as Arabic [23], Serbian, Tagalog, or Swahili, particularly when users apply spelling variation or obfuscation to evade filters. Code-switching and transliteration further degrade detection performance, as the model may miss common patterns or interpret them out of context [24, 25].
Cross-lingual models and prompt-tuned systems have shown moderate gains, especially when fine-tuned on translated or augmented datasets. However, language imbalance persists, especially in online spaces with small but highly active non-English user communities. For platforms with global audiences, this introduces a serious equity issue: harmful content in English is far more likely to be caught than equivalent abuse in low-resource languages or regional dialects.
II-D Evaluating LLMs for Safety and Fairness in Moderation
The rise of general-purpose LLMs like GPT-4, Claude, and Gemini has led to growing interest in using them for content moderation. Some recent work shows that these models can outperform task-specific classifiers in zero-shot settings, especially on unseen data [13, 21]. They are also more flexible, capable of incorporating platform policies or contextual reasoning in their prompts. However, their moderation performance can be highly sensitive to prompt phrasing, target domain, and cultural framing.
More recent research has introduced sociocultural audit frameworks, including persona-based prompts [27] and synthetic demographic simulations, to evaluate fairness and blind spots. These studies highlight that model behavior often shifts depending on perceived user identity or topic domain, raising concerns about bias and overgeneralization [28]. While LLMs offer clear advantages in flexibility and scalability, their reliance on pretraining data and lack of transparency in decision-making still pose barriers to safe deployment in moderation workflows.
This study builds on these gaps by evaluating leading LLMs in a real-world moderation scenario: cyberbullying detection in emotionally charged YouTube comment sections, covering multiple languages and informal speech styles. It extends previous work by focusing not only on overall accuracy but on model behavior across sarcasm, cultural cues, and under-resourced content.
III Methodology
III-A Case Selection and Data Collection
Comments were retrieved via the YouTube Data API between 15 April 2024 and 15 May 2025.
The final corpus covers four public videos from distinct creators and content domains such as gaming, lifestyle, food vlog, and music. To verify that each video was indeed a cyber-bullying hotspot, we first ran a pilot crawl of 1 000 comments per candidate clip and retained only those whose pilot abuse rate exceeded 20%. From the four retained videos we then drew a uniform random sample, yielding 5 080 comments in total. Only public endpoints were accessed; private, removed, or shadow-banned comments are not included.
III-B Data Cleaning and Anonymization
Each comment underwent a two–step preprocessing pipeline. First, text was normalised (UTF-8 decoding, whitespace trimming, emoji and punctuation were preserved). Second, personally identifiable information including user names, real names, phone numbers, e-mail addresses, links, and explicit geo-markers was either deleted or replaced by neutral placeholders such as UserNameProtected. No other linguistic content was altered. This procedure leaves the semantic core of each message intact while mitigating re-identification risk, allowing the corpus to be shared for research without exposing private data.
III-C Data Labeling and Ground Truth
All 5 080 comments were independently annotated by two reviewers (the first author and a second annotator holding a B.Sc. in Software Engineering) using a binary rubric:
Harmful
Bullying, sustained harassment, severe personal insults, or death-related threats
Not harmful
Neutral, supportive, off-topic, or otherwise non-abusive remarks
Inter‐rater reliability was high, with 91.0% raw agreement and $\kappa=0.83$, which is considered substantial. [^1] Landis and Koch, 1977. Disagreements (9 % of instances) were adjudicated by discussion until consensus, and that label was written to the final ground-truth file. Labels were never exposed to the models.
Non-English comments, including those in Arabic and Indonesian, were labeled using a combination of manual translation, consultation with fluent speakers, and, when necessary, language translation tools. While consistency was a priority, minor errors in interpreting slang or cultural references are possible and are noted as a study limitation.
III-D Evaluated Models
Three widely used large language models were selected for evaluation:
  1. 1.
    OpenAI GPT-4.1
  2. 2.
    Google Gemini 1.5 Pro
  3. 3.
    Anthropic Claude 3 Opus (2024-02-29)
Each model was tasked with classifying the same 5 080 user-generated comments using identical prompt structures to ensure fairness and consistency across systems. These models were selected due to their wide adoption, advanced reasoning capabilities, and public accessibility through API services, making them representative of current best practices in AI content moderation.
III-E Prompt Design
All three models were queried with the identical, minimal instruction shown below. Temperature was fixed at 0.0 and the maximum response length at 5 tokens to enforce deterministic, one–token outputs.
Analyze the following comment and determine if it contains cyberbullying, harassment, or harmful language. Make sure to understand the context, tone and potential impact. Respond with 0 if it is not harmful and 1 if it is harmful. Comment is: <COMMENT_TEXT>
The prompt provides no label definitions beyond the binary mapping, compelling each model to infer an internal policy from the plain-language description while still accounting for nuance.
III-F Evaluation Metrics
Each model’s output was compared against the manually labeled ground truth using standard classification metrics:
  • True Positives (TP): Model correctly flagged a harmful comment
  • True Negatives (TN): Model correctly identified a non-harmful comment
  • False Positives (FP): Model incorrectly flagged a non-harmful comment
  • False Negatives (FN): Model failed to flag a harmful comment
From this, the following performance metrics were calculated:
Precision measures the proportion of comments that the model correctly identified as harmful out of all the comments it flagged as harmful. In other words, it reflects how accurate the model is when it predicts that a comment is harmful. It is calculated using Equation 1.
$\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$ (1)
Recall looks at all the harmful comments in the dataset and measures how many the model correctly identified. A higher recall means the model caught more of the actual harmful content. The corresponding formula is shown in Equation 2.
$\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ (2)
F1 Score provides a balance between precision and recall. It is especially useful when both false positives and false negatives carry equal importance. The formula is defined in Equation 3.
[…]
Accuracy shows the overall percentage of correct predictions, covering both harmful and non-harmful cases. While it gives a general sense of performance, it may be less informative in imbalanced datasets. The formula appears in Equation 4.
$\text{Accuracy}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}$ (4)
Together, these metrics offer a well-rounded view of the models’ moderation performance across different types of classification errors.
Macro-averaged (class-balanced) scores were also calculated with the 26 % / 74 % split and yielded the same model ranking, so the detailed numbers are omitted for brevity.
III-G Manual Review and Model Comparison
Quantitative scores alone do not reveal why a system succeeds or fails, so we carried out a post-hoc qualitative review. Two annotators independently inspected a stratified sample of 200 disagreements, consisting of 50 false positives, 50 false negatives, and 100 edge-case ties, drawn in equal proportion from all three models. For each comment we discussed (i) whether the human ground truth should stand and (ii) what linguistic cues might have misled the model (sarcasm, coded slurs, emoji, topic drift, and so forth). Notes from this exercise were grouped into recurring error themes that inform the Discussion section. The procedure does not alter any numeric results but clarifies how each system handles nuance, context, and cultural references.
IV Results
IV-A Dataset Composition
The evaluation dataset consisted of 5 080 comments, including 1 334 harmful and 3 746 non-harmful instances. Harmful content made up about 26.3 percent of the total dataset, indicating a moderate class imbalance that could influence model metrics like precision and recall. Table I provides the class breakdown.
TABLE I: Dataset Distribution
Class Count Percentage
Harmful 1 334 26.3%
Non-harmful 3 746 73.7%
Total 5 080 100%
IV-B Model-Level Classification Performance
Each model was evaluated using raw classification outcomes and derived performance metrics. Table II presents the counts for true positives (TP), true negatives (TN), false positives (FP), and false negatives (FN) for each model.
TABLE II: Prediction Counts for Each Model
Model TP TN FP FN
GPT 1 122 3 603 143 212
Gemini 1 167 3 392 354 167
Claude 961 3 663 83 373
All three models showed strong ability to correctly identify non-harmful comments, with true negative counts above 3 300. Gemini identified the highest number of harmful comments, with 1 167 true positives, but also had the highest number of false positives, mistakenly labeling 354 non-harmful comments as harmful. GPT had a solid performance overall, producing 1 122 true positives and only 143 false positives. Claude was more selective, correctly identifying 961 harmful comments while minimizing false positives to 83.
From these outcomes, standard classification metrics were calculated: accuracy, precision, recall, F1 score, and false positive rate. These provide a clearer picture of each model’s behavior in detecting harmful content. Table III displays the results.
TABLE III: Evaluation Metrics Based on Prediction Outcomes
Model Accuracy Precision Recall F1 Score FPR
GPT 0.930 0.887 0.841 0.863 0.038
Gemini 0.897 0.767 0.875 0.818 0.095
Claude 0.910 0.920 0.720 0.808 0.022
GPT achieved the most balanced performance across the board. It reached an F1 score of 0.863, with high values for both precision at 0.887 and recall at 0.841. Claude stood out for its high precision, achieving 0.920, and maintained the lowest false positive rate at 0.022. However, this conservative stance came at the cost of recall, which was limited to 0.72. Gemini prioritized identifying as many harmful comments as possible, leading to the highest recall of 0.875, but its precision dropped to 0.767 and its false positive rate increased to 0.095.
IV-C Observations and Comparisons
The models displayed clear differences in their handling of harmful content. GPT offered a strong balance, combining effective detection with a relatively low error rate. This made it suitable for environments where both safety and user experience matter. Claude’s high precision and low false positive rate reflect a more cautious approach, making it better for contexts where false accusations must be minimized. Gemini, with the highest recall, is more aggressive, potentially fitting platforms that prioritize safety even at the cost of occasional over-flagging.
Each model reflects a distinct moderation philosophy. GPT is consistent and balanced, Claude is precise and conservative, and Gemini is proactive and broad-reaching. The choice of model should depend on the platform’s goals and tolerance for errors in either direction.
IV-D Visual Comparison
Figure 1 compares the models using precision, recall, and F1 score. These metrics highlight how each model prioritizes different aspects of content moderation.
Figure 1: Comparison of precision, recall, and F1 score across evaluated models.
The visual comparison supports earlier findings. GPT maintained a steady balance between precision and recall. Claude achieved top precision but lagged in recall. Gemini led in recall, though with lower precision. These patterns underline that no single model is best across all metrics. Instead, the optimal choice depends on what kind of moderation a platform values most. A platform that can tolerate more false positives may prefer Gemini, while one focused on minimizing moderation mistakes may lean toward Claude. In higher-risk environments, a combination of models could help match the tone, topic, and user profile of each comment more effectively.
V Discussion
V-A False Positives and Surface-Level Flagging
To better understand model overflagging, we analyzed cases where AI models marked comments as harmful while human reviewers did not. These false positives were grouped by model agreement patterns to highlight shared tendencies and specific weaknesses. While most flagged content reflected strong language or emotional tone, it often lacked any abusive intent. Below are four common patterns observed.
V-A1 All Three Models vs. Human: Shared Overflagging Patterns
In a small number of cases, only 12 out of 5 080, all three models (GPT, Claude, and Gemini) identified comments as harmful, while human reviewers did not. These examples typically involved informal phrasing, sarcasm, or emotionally expressive language. For instance, “Y’all are so toxic” was likely flagged due to tone, despite not targeting a specific individual. Another example, “Omg…321lbs! That’s so sad,” lacked context and may have been flagged merely for mentioning weight. These instances illustrate how all three models tend to rely on surface-level cues, such as slang or emotional keywords, without fully assessing intent or context.
V-A2 GPT: Over-Flagging Emotional/Mental-Health Language
GPT’s false positives often involved emotionally charged or mental health-related phrasing that human reviewers did not consider harmful. These included expressions of concern, frustration, or informal critique that lacked abusive intent. GPT frequently flagged comments referencing emotional breakdowns, support for recovery, or dramatic language about personal change. While this may reflect a cautious stance toward sensitive topics, it also suggests an overreliance on trigger phrases without accounting for context or tone.
V-A3 Gemini: Over-Flagging Concern and Instability
Gemini frequently flagged comments that referenced mental health, personal change, or emotional concern, even when human reviewers found no clear harm. Many of these comments offered support, expressed worry, or questioned the authenticity of a user’s behavior in dramatic or sarcastic terms. Phrases like “please take care of yourself” or “you can’t keep scaring everyone like this” were often interpreted by Gemini as harmful. This suggests a tendency to overflag emotionally sensitive or speculative language, particularly when it touches on perceived instability or trauma.
V-A4 Claude: Over-Flagging Sarcasm and Hyperbole
Claude’s false positives often involved sarcastic remarks, informal roasts, or exaggerated critiques that lacked targeted hostility. Many flagged comments included casual profanity, meme references, or expressive language, for example, jokes about appearance, pop culture, or emotional overreactions. In several cases, Claude appeared to react to tone or strong phrasing rather than the presence of actual harm. This suggests that Claude may place greater weight on civility and politeness, leading it to flag socially edgy but benign content more frequently than human reviewers.
These patterns align with Table III: Claude’s high precision shows caution, Gemini’s lower precision reflects aggressiveness, and GPT sits between.
V-B False Negatives: Sarcasm, Subtext, and Missed Harm
V-B1 All Models vs. Human: Missed Sarcasm, Mockery, and Coded Harassment
Several harmful comments were missed by all three models, highlighting common blind spots around sarcasm, ridicule, and indirect hostility. Many of these remarks used emojis, mock praise, or exaggerated phrasing to insult, humiliate, or incite collective disdain. Examples included references to “clown emojis,” indirect threats like “he feeds off his haters, just ignore him to make him fall off,” or comparisons to breakdowns and mental illness framed as jokes.
V-B2 GPT: Missed Hostility Behind Humor
GPT missed a substantial number of harmful comments flagged by human reviewers due to sarcasm, coded ridicule, or indirect hostility. Phrases like “every time I see you in the hospital I smile” conveyed sustained mockery and dismissiveness toward someone’s well-being. These misses suggest that GPT may rely too heavily on surface-level tone and literal phrasing.
V-B3 Gemini: Missed Harm in Public Shaming
Gemini failed to flag a range of comments often involving mockery or aggression wrapped in sarcasm or cultural shorthand. Some escalated into public shaming, such as calling for mass unsubscribing or questioning the person’s sanity. Gemini’s tendency to underreact suggests it may struggle with patterns of collective ridicule.
V-B4 Claude: Missed Sarcasm and Faux Concern
Claude frequently overlooked comments like “Send this man back to Arkham,” which ridicule the creator through implication. It also failed to detect hostile rants accusing the creator of deception or instability, showing a reluctance to flag comments unless explicitly hostile.
V-C Risks of Missed Moderation in Sensitive Cases
False negatives are not equal in impact. While a single missed insult may seem minor, its harm can compound over time, especially in emotionally vulnerable environments. Several missed comments in this study targeted known vulnerabilities, mocked prior hospitalization, or questioned a creator’s mental health. These remarks, though subtle or sarcastic, reinforce harmful narratives and may discourage affected users from seeking help or participating in the platform.
More troubling were coordinated suggestions to ”unfollow so he falls off” or ”ignore him until he cracks again.” These comments reflect collective hostility rather than isolated aggression. When such messages go unflagged, they enable mob harassment campaigns that can lead to reputational harm, social withdrawal, or emotional distress for creators. In several cases, these remarks appeared alongside strings of other taunts, suggesting that their impact is not just in their wording but in their cumulative pressure.
This highlights a key limitation of single-comment moderation: by treating each comment in isolation, models miss larger patterns of piling-on or manipulation. Moderation systems must evolve to track conversation history, detect repeated targeting, and recognize coded language that signals coordinated behavior. Especially in cases involving mental health, even a few missed comments can shift the tone of a thread and undermine user safety. Future systems should include tools for temporal tracking, conversational memory, and risk-tiered escalation to reduce the long-term impact of missed moderation.
V-D Humor, Satire, and Ambiguity in Model Moderation
Another common source of disagreement between human reviewers and LLMs involved ambiguous comments that used humor, satire, or double meaning. These false positives often lacked direct hostility but contained emotionally charged phrasing, exaggerated tone, or social critique. Several examples reveal how models misinterpret stylistic or cultural cues.
One such comment was ”Can’t sing for nothing!!!!”, flagged by GPT despite being a common form of exaggerated critique in entertainment contexts. Similarly, ”2:33 too much jiggling” was flagged by Gemini, likely due to its focus on physical attributes, though the intent may have been commentary rather than harassment.
Comments referencing public figures or using layered sarcasm were also commonly misclassified. For instance, ”It sounds like UserNameProtected Cobain’s remembrance video” was flagged by GPT, Gemini, and Claude, though human reviewers judged it to be a stylistic comparison rather than an attack. Another flagged comment, ”You just gonna act like nothing happened?…” shows how models can mistake rhetorical or skeptical questions for aggression, especially when stripped of conversational context.
Claude and Gemini in particular flagged remarks like ”why does he look malnourished” or ”UserNameProtected ain’t even real” that are commonly used in meme culture or satire. These comments, while edgy, were not interpreted as harmful by human reviewers. However, the models appeared sensitive to tone and phrasing that hinted at ridicule or disbelief.
These cases underscore the limits of current models in interpreting intent, especially when humor blurs the line between critique and cruelty. Without access to conversational history or platform-specific norms, LLMs often err on the side of caution, flagging emotionally charged or socially playful content as harmful.
For moderation systems, this presents a difficult trade-off. Over-flagging benign humor risks alienating users and undermining trust in automated systems, while under-flagging allows veiled insults or passive aggression to persist. Future moderation pipelines may benefit from humor-aware classifiers or community-tuned thresholds that distinguish cultural satire from actual abuse.
VI Future Work
The study surfaces several gaps that invite deeper investigation and practical follow-up.
VI-A Thread-Level and Multi-modal Context
All three models evaluated a single comment in isolation. Future work should supply surrounding conversation threads, video transcripts, and visual cues so that the model can gauge sarcasm, piling-on, or context that flips a seemingly benign sentence into harassment.
VI-B Implicit and Coded Abuse
Many errors stemmed from sarcasm, memes, or coded references. A dedicated corpus that labels irony, metaphor, and insider jokes would let researchers test whether prompt engineering or specialised pre-training improves detection of covert hostility.
VI-C Ensemble and Risk-Tiered Pipelines
No single model led across every metric. A practical pipeline could run a high-recall model first, then pass flagged items through a high-precision filter before escalation to human review. Measuring cost, latency, and user impact of such cascades remains an open task.
VI-D User-Facing Feedback and Digital Citizenship
Flagging comments is only one lever. Interventions that nudge commenters toward civility, or that explain why language is harmful, might reduce abuse at the source. Controlled trials comparing silent removal, warning prompts, and educational pop-ups would clarify which approach changes behaviour most effectively.
VI-E Long-Term Outcomes and Mental-Health Impact
Finally, future studies should trace whether improved moderation actually lowers anxiety, self-harm ideation, or community churn. Partnering with mental-health researchers and using de-identified longitudinal data can turn model gains into real-world wellbeing.
VII Conclusion
This paper benchmarks OpenAI GPT-4.1, Google Gemini 1.5 Pro, and Anthropic Claude 3 Opus on 5 080 YouTube comments sampled from high-abuse threads in gaming, lifestyle, food vlog, and music channels. Each model was evaluated with the same prompt under deterministic settings. Gemini identified the largest share of harmful content with recall 0.875 but its precision fell to 0.767 due to frequent false positives. Claude achieved the highest precision 0.920 and the lowest false-positive rate 0.022 yet its recall slipped to 0.720. GPT delivered the best overall balance with an F1 score 0.863, precision 0.887, and recall 0.841. Qualitative analysis showed that sarcasm, coded insults, and mixed-language slang remain persistent blind spots for all three models. These findings make clear that no single system meets every moderation requirement. Practical pipelines should combine complementary models, incorporate conversational context, and fine-tune for under-represented languages and subtle forms of abuse. The fully de-identified dataset and detailed prompts are publicly released to support reproducible research and foster further improvements in automated content moderation.
VIII Limitations
This study has several limitations:
  1. 1.
    Zero-shot prompts. All models were evaluated using identical zero-shot prompts without model-specific fine-tuning or prompt optimization. This setup prioritizes comparability but may underrepresent each model’s best-case performance.
  2. 2.
    Comment-level context. Each comment was moderated in isolation, without access to prior messages or conversation history. This limits the ability to detect context-dependent abuse, sarcasm, or coordinated behavior across threads.
  3. 3.
    Cultural bias in labeling. Human annotations were provided by two reviewers with similar linguistic and cultural backgrounds, which may have introduced bias in interpreting sarcasm, slang, or ambiguous phrasing.
  4. 4.
    Limited language scope. Although the dataset includes occasional comments in Arabic or Indonesian, it was primarily designed for English-language analysis. As such, results should not be generalized to multilingual performance.
Ethical Considerations
This study uses only publicly available YouTube comments. All user identifiers and any personally identifying details were removed or replaced with neutral placeholders before analysis, so no individual can be re-identified. Because the data are de-identified and pose minimal risk, no formal IRB review was required.
Data Availability
The full, de-identified comment corpus, together with the human labels and model predictions, is openly available at https://github.com/Ammce/papers/tree/main/llm-cyberbullying-moderation%20. To minimise re-identification risk, user names, links, and YouTube comment IDs have been removed. The list of source‐video IDs can be shared with editors or qualified researchers under a non-disclosure agreement. Exact model prompts and API parameters used in this study are provided in the same repository under prompts/.